OpenStudy (anonymous):

simplify 3x/14x^2 x 7x^2/15x

OpenStudy (anonymous):

Cross multiply and this is how it looks: \[\frac{ 3x }{ 14x^{2} } \times \frac{ 7x^{2} }{ 15x }\] Then: \[\frac{ 1 }{ 2x^{2} } \times \frac{ 1 }{ 5x }\]

OpenStudy (anonymous):

@AnyTipzWouldHelp would I have to cancel out exponents? & thank you so much

OpenStudy (anonymous):

Your final answer is: \[\frac{ 1 }{ 10 }\] I did all the other work for you since it's easy and it's just busy work :)